# Erwin: A Tree-based Hierarchical Transformer for Large-scale Physical Systems

### v1

@@ 5.2 Molecular dynamics
Molecular dynamics (MD) is essential for understanding physical and biological systems at the atomic level but remains computationally expensive even with neural network potentials due to all-atom force calculations and femtosecond timesteps required to maintain stability and accuracy. Fu et al. (2022) suggested accelerating MD simulation through coarse-grained dynamics with MPNN. In this experiment, we take a different approach and instead operate on the original representation but improve the runtime by employing our hardware-efficient model. Therefore, the question we ask is how much we can accelerate a simulation w.r.t. an MPNN without compromising the performance.
@@
The dataset consists of single-chain coarse-grained polymers (Webb et al., 2020; Fu et al., 2022) simulated using MD. Each system includes 4 types of coarse-grained beads interacting through bond, angle, dihedral, and non-bonded potentials. The training set consists of polymers with the repeated pattern of the beads while the polymers in the test set are constructed by randomly sampling sequences of the beads thus introducing a challenging distribution shift. The training set contains $100$ short trajectories ($50k$ $\tau$), while the test set contains $40$ trajectories that are $100$ times longer. Each polymer chain contains approximately $890$ beads on average.
@@
We follow the experimental setup from Fu et al. (2022). The model takes as input a polymer chain of $N$ coarse-grained beads. Each bead has a specific weight and is associated with the history $\{\dot{\mathbf{x}}_{t-16\Delta t},...,\dot{\mathbf{x}}_{t-\Delta t}\}$ of (normalized) velocities from $16$ previous timesteps at intervals of $\Delta t=5\tau$. The model predicts the mean $\mathbf{\mu}_{t}\in\mathbb{R}^{N\times 3}$ and variance $\mathbf{\sigma}_{t}^{2}\in\mathbb{R}_{+}^{N\times 3}$ of (normalized) acceleration for each bead, assuming a normal distribution. We train using negative log-likelihood loss
$\mathcal{L}=-\log\mathcal{N}(\hat{\ddot{\mathbf{x}}}_{t}|\mathbf{\mu}_{t},\mathbf{\sigma}_{t}^{2})\vspace{-2pt}$
between predicted and ground truth accelerations computed from the ground truth trajectories.
@@
Table 2: Ablation study (MD task, Erwin-S) on the architectural choices: using MPNN in embedding, RPE (see Eq. 9) and cross-ball connection via rotating trees. Runtime includes building trees.
Test Loss Runtime, ms
w/o $0.738$ $25.15$
   $+$ MPNN $0.720$ $26.07$
   $+$ RPE $0.715$ $26.05$
   $+$ rotating tree $0.712$ $26.73$
@@ Results
The results are given in Fig. 7. As baselines, we use MPNN (Gilmer et al., 2017) as well as two hardware-efficient architectures: PointNet++ (Qi et al., 2017) and PointTransformer v3 (Wu et al., 2024). Notably, model choice has minimal impact on performance, potentially due to the absence of long-range interactions as the CG beads do not carry any charge. Furthermore, it is sufficient to only learn local bonded interactions. There is, however, a considerable improvement in runtime for Erwin ($1.7-2.5$ times depending on the size), which is only matched by smaller MPNN or PointNet++, both having significantly higher test loss.
@@ 5.3 Turbulent fluid dynamics
In the last experiment, we demonstrate the expressivity of our model by simulating turbulent fluid dynamics. The problem is notoriously challenging due to multiple factors: the inherently nonlinear behaviour of fluids, the multiscale and chaotic nature of turbulence, and the presence of long-range dependencies. Moreover, the geometry of the simulation domain and the presence of objects introduce complex boundary conditions thus adding another layer of complexity.
@@
We use EAGLE (Janny et al., 2023), a large-scale benchmark of unsteady fluid dynamics. Each simulation includes a flow source (drone) that moves in 2D environments with different boundary geometries producing airflow. The time evolution of velocity and pressure fields is recorded along with dynamically adapting meshes. The dataset contains $600$ different geometries of $3$ types, with approximately $1.1$ million 2D meshes averaging $3388$ nodes each. The total dataset includes $1184$ simulations with $990$ time steps per simulation. The dataset is split with 80% for training and 10% each for validation and testing.
@@
We follow the original experimental setup of the benchmark. The input is the velocity $V\in\mathbb{R}^{N\times 2}$ and pressure $P\in\mathbb{R}^{N\times 2}$ fields evaluated at every node of the mesh in the time step $t$ along with the type of the node. The task is to predict the state of the system at the next time step $t+1$. The training is done by predicting a trajectory of states of length $5$ and optimizing the loss
@@ Results
For comparison, we include the baselines from the original benchmark: MeshGraphNet (MGN; Pfaff et al., 2021), GAT (Velickovic et al., 2018), DilResNet (DRN; Stachenfeld et al., 2021) and EAGLE (Janny et al., 2023) [^6] We additionally trained UPT (Alkin et al., 2024a), but were not able to obtain competitive results in our initial experiments.. The first two baselines are based on message-passing, while DilResNet operates on regular grids hence employing interpolation for non-uniform meshes. EAGLE uses message-passing to pool the mesh to a coarser representation with a fixed number of clusters, on which attention is then computed. The quantitative results are given in Table 3 and unrolling trajectories are shown in Fig. 8 and in Appendix B. Erwin demonstrates strong results on the benchmark and outperforms every baseline, performing especially well at predicting pressure. In terms of inference time and memory consumption, Erwin achieves substantial gains over EAGLE, being $3$ times faster and using $8$ times less memory.
Table 3: RMSE on velocity V and pressure P fields across different prediction horizons (mean $\pm$ std over 5 runs). Inference runtime and memory use computed for a batch of $8$, $3500$ nodes on average.
@@
5.4 Ablation study
We also conducted an ablation study to examine the effect of increasing ball sizes on the model’s performance in the cosmology experiment, see Table 1. Given the presence of long-range interactions in the data, larger window sizes (and thus receptive fields) improve model performance, albeit at the cost of increased computational runtime. Our architectural ablation study on the MD task (Table 2) reveals that using MPNN at the embedding step produces substantial improvements, likely due to its effectiveness in learning local interactions.
6 Conclusion
We present Erwin, a hierarchical transformer that uses ball tree partitioning to process large-scale physical systems with linear complexity. Erwin achieves state-of-the-art performance on both the cosmology benchmark (Balla et al., 2024) and the EAGLE dataset (Janny et al., 2023), demonstrating its effectiveness across diverse physical domains. The efficiency of Erwin makes it a suitable candidate for any tasks that require modeling large particle systems, such as tasks in computational chemistry (Fu et al., 2024) or diffusion-based molecular dynamics (Jing et al., 2024).
Limitations and Future Work
Erwin was developed by jointly optimizing for expressivity and runtime. As a result, certain architectural decisions are not optimal with respect to memory usage. In particular, we use a distance-based attention bias (see Eq. 10), for which both computational and memory requirements grow quadratically with the ball size. Developing alternative ways of introducing geometric information into attention computation could reduce these requirements. Finally, Erwin is neither permutation nor rotation equivariant, although rotation equivariance can be incorporated without compromising scalability. One possible approach is to use geometric algebra transformers (Brehmer et al., 2023) and omit the proposed cross-ball connections, as they rely on invariance-breaking tree building.
Acknowledgement
We are grateful to Evgenii Egorov and Ana Lučić for their feedback and inspiration. This research was supported by Microsoft Research AI4Science.
@@ Appendix A Implementation details
The algorithm used for constructing ball trees (Pedregosa et al., 2012) can be found in Alg. 1. Note that this implementation is not rotationally equivariant as it relies on choosing the dimension of the greatest spread which in turn depends on the original orientation. Examples of ball trees built in our experiments are shown in Fig. 9.
@@
where $\mathbf{h}_{i}\in\mathbb{R}^{H}$ is a feature vector of $v_{i}$, $\mathcal{N}(i)$ denotes the neighborhood of $v_{i}$. The motivation for using an MPNN is to incorporate local neighborhood information into the model. Theoretically, attention should be able to capture it as well; however, this might require substantially increasing feature dimension and the number of attention heads, which would be prohibitively expensive for a large number of nodes in the original level of a ball tree.
In our experiments, we consistently maintain the size of MLPe and MLPh small ($H\leq 32$) such that embedding accounts for less than $5$% of total runtime.
Appendix B Extended experiments
B.1 Turbulent fluid dynamics
We provide additional exemplary rollouts of Erwin for both velocity (Fig. 10) and pressure (Fig. 11) fields.
B.2 Airflow pressure modeling
Dataset
We use the ShapeNet-Car dataset generated by Umetani & Bickel (2018) and preprocessed by Alkin et al. (2024a). It consists of $889$ car models, each car being represented by $3586$ surface points in $3$D space. Airflow was simulated around each car for $10$s (Reynolds number $Re=5\times 10^{6}$) and averaged over the last $4$s to obtain pressure values at each point. The dataset is randomly split into $700$ training and $189$ test samples.
Task
Given surface points, the task is to predict the value of pressure $P\in\mathbb{R}^{N\times 1}$ at each point in $X^{N\times 3}$. The training is done by optimizing the mean squared error loss between predicted and ground truth pressures.
Table 4: Test MSE for ShapeNet-Car pressure prediction. [U] and [R] indicate U-Net- and ResNet-like structures respectively.
Model MSE, ×$\times 10^{-2}$
U-Net $6.13$
FNO $4.04$
GINO $2.34$
UPT $2.31$
PointTransformer v3 [U] $1.78$
Erwin (Ours) [U] $1.42$
PointTransformer v3 [R] $\mathbf{0.92}$
Erwin (Ours) [R] $\mathbf{0.92}$
Results
The results are given in Table 4. We evaluate PointTransformer v3 and use the baseline results obtained by Alkin et al. (2024a) for U-Net (Ronneberger et al., 2015), FNO (Li et al., 2021), GINO (Li et al., 2023), and UPT (Alkin et al., 2024a). Both Erwin and PointTransformer v3 achieve significantly lower test MSE compared to other models, which can be attributed to their ability to capture fine geometric details by operating directly on the original point cloud. In comparison, other approaches introduce information loss through compression - UPT encodes the mesh into a latent space representation, while the remaining baselines interpolate the geometry onto regular grids and back. Moreover, in our experiments, ResNet-like configurations that did not include any coarsening performed dramatically better than the ones following the U-Net structure. Overall, this result highlights the potential for Erwin to be used as a scalable neural operator (Wang et al., 2024).
@@ Results
Appendix C Experimental details
In this section, we provide experimental details regarding hyperparameter choice and optimization. All experiments were conducted on a single NVIDIA RTX A6000. All models were trained using the AdamW optimizer (Loshchilov & Hutter, 2019) with weight decay $0.01$ and a cosine decay schedule. The learning rate was tuned in the range $10^{-4}$ to $10^{-3}$ to minimize loss on the respective validation sets.

### v2

@@ 5.2 Molecular dynamics
Molecular dynamics (MD) is essential for understanding physical and biological systems at the atomic level but remains computationally expensive even with neural network potentials due to all-atom force calculations and femtosecond timesteps required to maintain stability and accuracy. Fu et al. (2022) suggested accelerating MD simulation through coarse-grained dynamics with an MPNN. In this experiment, we take a different approach and instead operate on the original representation but improve the runtime by employing our hardware-efficient model. Therefore, the question we ask is how much we can accelerate a simulation w.r.t. an MPNN without compromising the performance.
@@
The dataset consists of single-chain coarse-grained polymers (Webb et al., 2020; Fu et al., 2022) simulated using MD. Each system includes 4 types of coarse-grained beads interacting through bond, angle, dihedral, and non-bonded potentials. The training set consists of polymers with repeated bead patterns while the test set polymers are constructed by randomly sampling bead sequences, thus introducing a challenging distribution shift. The training set contains $100$ short trajectories (50k $\tau$), while the test set contains 40 trajectories that are 100 times longer.
@@
We follow the experimental setup from Fu et al. (2022). The model takes as input a polymer chain of $N$ coarse-grained beads. Each bead has a specific weight and is associated with the history $\{\dot{\mathbf{x}}_{t-16\Delta t},...,\dot{\mathbf{x}}_{t-\Delta t}\}$ of (normalized) velocities from 16 previous timesteps at intervals of $\Delta t=5\tau$. The model predicts the mean $\mathbf{\mu}_{t}\in\mathbb{R}^{N\times 3}$ and variance $\mathbf{\sigma}_{t}^{2}\in\mathbb{R}_{+}^{N\times 3}$ of (normalized) acceleration for each bead, assuming a normal distribution. We train using the negative log-likelihood loss between predicted and ground truth accelerations computed from the ground truth trajectories.
@@
Table 3: Ablation study on architectural choices: using MPNN in embedding, RPE, and cross-ball connection via rotating trees.
Model Test Loss
MD ShapeNet-Car
w/o $0.738$ $30.39$
   $+$ MPNN $0.720$ $30.49$
   $+$ RPE $0.715$ $30.02$
   $+$ rotating tree $0.712$ $15.85$
@@ Results
Table 4: RMSE on PDE benchmarks from Li et al. (2023b). Transformer-based models are taken as baselines from Luo et al. (2025).
Model RMSE
Elasticity Plasticity Airfoil Pipe
LNO (2024) 0.69 0.29 0.53 0.31
Galerkin (2021) 2.40 1.20 1.18 0.98
HT-Net (2022) / 3.33 0.65 0.59
OFormer (2023c) 1.83 0.17 1.83 1.68
GNOT (2023) 0.86 3.36 0.76 0.47
FactFormer (2023d) / 3.12 0.71 0.60
ONO (2024) 1.18 0.48 0.61 0.52
Transolver++ (2025) 0.52 0.11 0.48 0.27
Erwin (Ours) 0.34 0.10 2.57 0.61
Table 5: Test MSE for ShapeNet-Car pressure prediction. Baseline results are taken from Bleeker et al. (2025).
Model MSE
PointNet (2017) $43.36$
GINO (2023b) $35.24$
UPT (2024a) $31.66$
Transolver (2024a) $19.88$
GP-UPT (2025) 17.02
PTv3-S (2024b) $19.09\pm 0.67$
PTv3-M (2024b) $17.42\pm 0.38$
Erwin-S (Ours) $\mathbf{15.85\pm 0.19}$
Erwin-M (Ours) $\mathbf{15.43\pm 0.45}$
The results are given in Fig. 7. As baselines, we use MPNN (Gilmer et al., 2017) as well as two hardware-efficient architectures: PointNet++ (Qi et al., 2017) and PTv3 (Wu et al., 2024b). Notably, model choice has minimal impact on performance, potentially due to the absence of long-range interactions as the CG beads do not carry any charge. Furthermore, it is sufficient to only learn local bonded interactions. There is, however, a considerable improvement in runtime for Erwin (1.7-2.5 times depending on the size), which is only matched by a smaller MPNN or PointNet++, both having significantly higher test loss.
5.3 PDE benchmarks and airflow pressure
Deep learning models have emerged as surrogate solvers of partial differential equations (PDEs), learning to approximate solutions from data (Li et al., 2021; Wu et al., 2024a). While their advantages over traditional numerical methods remain unclear (McGreivy & Hakim, 2024), the task itself serves as a surrogate for large-scale applications like weather forecasting (Price et al., 2025) and fluid dynamics (Bleeker et al., 2025), where conventional solvers become computationally prohibitive. Furthermore, in this task, we are interested in the model’s ability to scale to large domains while capturing complex patterns in underlying physics.
Dataset
We benchmark on multiple datasets taken from Li et al. (2023a). Each dataset is defined either on point cloud (Elasticity) or structured mesh (Plasticity, Airfoil, Pipe). Additionally, we evaluate our model on airflow pressure modeling (Umetani & Bickel, 2018; Alkin et al., 2024a). It consists of 889 car models, each car represented by 3,586 surface points in 3D space. Airflow was simulated around each car for $10$s ($Re=5\times 10^{6}$) and averaged over the last $4$ s to obtain pressure values at each point.
Task
For the PDE benchmarks, we follow the pipeline from Wu et al. (2024a) and minimize the relative L2 error; see Appendix B for details. For ShapeNet-Car, the task is to estimate the surface pressure $\mathbf{Y}\in\mathbb{R}^{3388\times 1}$ given surface points $\mathbf{X}\in\mathbb{R}^{3388\times 3}$. We train by optimizing the mean squared error between predicted and ground truth pressure.
PDE benchmarks
The results are given in Table 4, where we compare against other transformer-based methods. Erwin achieves state-of-the-art performance on 2 out of 4 tasks. Interestingly, it dramatically underperforms on the Airfoil task, which indicates a failure mode. We speculate that this is related to a specific structure of the data - the density of the mesh decreases dramatically moving away from the center of mass. This means that points across different balls have considerably varying density, which poses a challenge.
ShapeNet-Car
See Table 5 for results. Both Erwin and PTv3 achieve significantly lower test MSE compared to other models [^6] The baseline results for the ShapeNet-Car task, except for PTv3, were taken from Bleeker et al. (2025).. We note that the best performing configuration of Erwin and PTv3 did not include any coarsening, thus operating directly on the original point cloud. This indicates that the task favors the ability of a model to capture fine geometric details. In comparison, other approaches introduce information loss through compression - UPT (Alkin et al., 2024a) and Transolver (Wu et al., 2024a) involve pooling to the latent space, while GINO (Li et al., 2023b) interpolates the geometry onto regular grids and back.
@@ 5.4 Turbulent fluid dynamics
In the last experiment, we demonstrate the expressivity of our model by simulating turbulent fluid dynamics. The problem is notoriously challenging due to multiple factors: the inherently nonlinear behavior of fluids, the multiscale and chaotic nature of turbulence, and the presence of long-range dependencies. Moreover, the geometry of the simulation domain and the presence of objects introduce complex boundary conditions, thus adding another layer of complexity.
@@
We use EAGLE (Janny et al., 2023), a large-scale benchmark of unsteady fluid dynamics. Each simulation includes a flow source (drone) that moves in 2D environments with different boundary geometries, producing airflow. The time evolution of velocity and pressure fields is recorded along with dynamically adapting meshes. The dataset contains 600 different geometries of 3 types, with approximately 1.1 million 2D meshes averaging 3,388 nodes each. The total dataset includes 1,184 simulations with 990 time steps per simulation. The dataset is split with 80% for training and 10% each for validation and testing.
@@
We follow the original experimental setup of the benchmark. The input is the velocity $V\in\mathbb{R}^{N\times 2}$ and pressure $P\in\mathbb{R}^{N\times 2}$ fields evaluated at every node of the mesh at time step $t$, along with the type of the node. The task is to predict the state of the system at the next time step $t+1$. The training is done by predicting a trajectory of states of length 5 and optimizing the loss
@@ Results
For comparison, we include the baselines from the original benchmark: MeshGraphNet (MGN; Pfaff et al., 2021), GAT (Velickovic et al., 2018), DilResNet (DRN; Stachenfeld et al., 2021) and EAGLE (Janny et al., 2023) [^7] We additionally trained UPT (Alkin et al., 2024a), but were not able to obtain competitive results in our initial experiments.. The first two baselines are based on message-passing, while DilResNet operates on regular grids, hence employing interpolation for non-uniform meshes. EAGLE uses message-passing to pool the mesh to a coarser representation with a fixed number of clusters, on which attention is then computed. The quantitative results are given in Table 6 and unrolling trajectories are shown in Fig. 89. Erwin demonstrates strong results on the benchmark and outperforms every baseline, performing especially well at predicting pressure. In terms of inference time and memory consumption, Erwin achieves substantial gains over EAGLE, being 3 times faster and using 8 times less memory.
5.5 Ablation study
We also conducted an ablation study to examine the effect of increasing ball sizes on the model’s performance in the cosmology experiment; see Table 2. Given the presence of long-range interactions in the data, larger window sizes (and thus receptive fields) improve model performance, albeit at the cost of increased computational runtime. Our architectural ablation study on the MD task (Table 3) reveals that using an MPNN at the embedding step produces substantial improvements, likely due to its effectiveness in learning local interactions. At the same time, that did not generalize to ShapeNet-Car, where the most decisive factor was including cross-ball interactions, highlighting the importance of capturing fine-grained details in this task.
6 Conclusion
We present Erwin, a hierarchical transformer that uses ball tree partitioning to process large-scale physical systems with linear complexity. Erwin achieves state-of-the-art performance in cosmology, turbulent fluid dynamics, and, partially, on standard PDE benchmarks, demonstrating its effectiveness across diverse physical domains. The efficiency of Erwin makes it a suitable candidate for tasks that require modeling large particle systems, such as computational chemistry (Fu et al., 2024) or diffusion-based molecular dynamics (Jing et al., 2024).
Limitations and Future Work
Table 6: RMSE on velocity V and pressure P fields across different prediction horizons (mean $\pm$ std over 5 runs). Inference runtime and memory use are computed for a batch of 8, avg. 3,500 nodes.
@@ 6 Conclusion
Erwin relies on cross-ball interaction and coarsening to capture long-range interactions. Both mechanisms have inherent limitations: the former requires multiple steps for signals to propagate between balls, while the latter sacrifices fine-grained detail. A promising approach to address these issues is adapting sparse attention methods such as Native Sparse Attention (Yuan et al., 2025). This framework aligns naturally with Erwin’s ball tree structure and would enable learning distant interactions while preserving full resolution. Finally, Erwin is neither permutation nor rotation equivariant, although rotation equivariance can be incorporated without compromising scalability, such as via Geometric Algebra Transformers (Brehmer et al., 2023) or Fast Euclidean Attention (Frank et al., 2024).
Acknowledgements
We are grateful to Evgenii Egorov and Ana Lučić for their feedback and inspiration. MZ acknowledges support from Microsoft Research AI4Science. JWvdM acknowledges support from the European Union Horizon Framework Programme (Grant agreement ID: 101120237).
Impact statement
The broader implications of our work are primarily in moderate- to large-scale scientific applications, such as molecular dynamics or computational fluid dynamics. We believe that efficient and expressive architectures like Erwin could become a foundation for resource-intensive deep learning frameworks and therefore help in better understanding the physical systems governing our world.
@@ Appendix A Implementation details
The algorithm used for constructing ball trees (Pedregosa et al., 2012) can be found in Alg. 1. Note that this implementation is not rotationally equivariant as it relies on choosing the dimension of the greatest spread, which in turn depends on the original orientation. Examples of ball trees built in our experiments are shown in Fig. 9.
@@
where $\mathbf{h}_{i}\in\mathbb{R}^{H}$ is a feature vector of $v_{i}$, and $\mathcal{N}(i)$ denotes the neighborhood of $v_{i}$. The motivation for using an MPNN is to incorporate local neighborhood information into the model. Theoretically, attention should be able to capture this information as well; however, this might require substantially increasing the feature dimension and the number of attention heads, which would be prohibitively expensive for a large number of nodes in the original level of a ball tree. In our experiments, we consistently maintain the size of MLPe and MLPh to be small ($H\leq 32$) such that embedding accounts for less than $5$% of total runtime.
@@
Appendix B Experimental details
In this section, we provide comprehensive experimental details regarding hardware specifications, hyperparameter choices, and optimization procedures.
B.1 Hardware and Software
All experiments were conducted on a single NVIDIA RTX A6000 GPU with 48GB memory and 16 AMD EPYC™ 7543 CPUs. Erwin and all baselines except those for cosmology were implemented in PyTorch 2.6. For the cosmology benchmark, SEGNN, NequIP, and MPNN baselines were implemented in JAX as provided by the original benchmark repository. Training times for Erwin varied by task: cosmology (5-10 minutes depending on training set size), molecular dynamics (2-4 hours depending on model size), PDE benchmarks (8 hours for Elasticity, 48 hours for others), ShapeNet-Car (2 hours), and EAGLE (48 hours).
B.2 Training Details
All models were trained using the AdamW optimizer (Loshchilov & Hutter, 2019) with weight decay $10^{-5}$. The learning rate was tuned in the range $10^{-4}$ to $10^{-3}$ to minimize loss on the respective validation sets with cosine decay to $10^{-7}$. Gradient clipping by norm with value 1.0 was applied across all experiments. Early stopping was used only for ShapeNet-Car and molecular dynamics tasks, while all other models were trained until convergence. In every experiment, we normalize inputs to the model. Hyperparameter optimization was performed using grid search with single trials.
B.3 Dataset Splits and Evaluation
Dataset splits followed the original benchmarks:
  • Cosmology: Training set varied from 64 to 8192 examples, with validation and test sets of 512 examples each
  • Molecular Dynamics: 100 short trajectories for training, 40 long trajectories for testing
  • PDE Benchmarks: 1000 training / 200 test examples (except Plasticity: 900/80)
  • ShapeNet-Car: 700 training / 189 test examples
  • EAGLE: 1184 trajectories with 80%/10%/10% split
For statistical significance, we ran each experiment 5 times and report mean and standard deviation for cosmology, molecular dynamics, ShapeNet-Car, and EAGLE. PDE experiments were run once due to computational constraints.
B.4 Evaluation Metrics
For RMSE computation, we use the relative L2 error: $\text{RMSE}:=\|f(x)-y\|/\|y\|$. For molecular dynamics, we randomly sample 16 history points from each trajectory and predict one future point. Acceleration is predicted by the neural network based on history and compared against ground truth computed using forward differences. Model performance is evaluated using negative log-likelihood loss between predicted and ground truth accelerations.
B.5 Baseline Implementations
All baseline results were taken from official implementations or reported values:
B.6 Computational Efficiency
Table 7 shows the inference time breakdown for Erwin on NVIDIA RTX A6000 with batch size 16. Ball tree construction (Table 8) consistently accounts for less than 5% of total runtime, demonstrating the efficiency of our optimized implementation compared to standard libraries.
Table 7: Erwin runtime with torch.compile
Nodes per batch Runtime (ms)
16 $\times$ 2048 4096 8192 16384
Fwd 17.3 31.6 79.7 189
Fwd + Bwd 26.4 45.4 114 232
Table 8: Ball tree construction on 16 AMD EPYC™ 7543 CPUs.
Nodes per batch, Runtime (ms)
16 $\times$ 2048 4096 8192 16384
sklearn + joblib 16.3 21.2 24.1 44.0
Ours 0.73 1.54 3.26 6.98
Speed-up 22.3× 13.8× 7.4× 6.3×
B.7 Further details per experiment